I don't know if I'm posting this in the right area but could you please help me. I need to find the distance between the points (-4,-6) and (-8,2). Simplify your answer.
Found 2 solutions by swctaker, venugopalramana:
Answer by swctaker(2)

(-4-(-8))/(-6-2)= 4/-8 = -1/2

Answer by venugopalramana(3286)

I don't know if I'm posting this in the right area but could you please help me. I need to find the distance between the points (-4,-6) and (-8,2). Simplify your answer.
DISTANCE BETWEEN 2 POINTS (X1,Y1) AND (X2,Y2) IS GIVEN BY
D=SQRT.{(X1-X2)^2+(Y1-Y2)^2}={(-4+8)^2+(-6-2)^2}^0.5=(16+64)^0.5=80^0.5=4*SQRT(5)
=8.944